OpenStudy (anonymous):

(cosx/tanx)(1-sec^2(x))

OpenStudy (anonymous):

Another one with trig identities?

OpenStudy (anonymous):

I am really bad a these lol

OpenStudy (anonymous):

This is another on that uses the fact that you gotta simplify tan first and now we're gonna use another new one: \[\Huge 1+\cot^2(x)=\sec^2(x)\]

OpenStudy (anonymous):

so 1-sec^2(x) would become 1-1+cot^2(x)--> cot^(2)x?

OpenStudy (anonymous):

No sorry.

OpenStudy (anonymous):

\[\Huge 1+\tan^2(x)=\sec^2(x)\]

OpenStudy (anonymous):

so 1-1+tan^2(x)--> tan^2x?

OpenStudy (anonymous):

Okay, so I plugged in what I just got and my end result is sinx, is that right?

OpenStudy (anonymous):

I found on a website that tan^2(x)+1=sec^2(x), does the position of the "1" matter?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Well on the one side of the equation, no

OpenStudy (anonymous):

Like you can have 1+tan^2(x)=sec^2(x) or tan^2(x)+1=sec^2(x) but you can't have somehting like tan^2(x)=sec^2(x)+1 or tan^2(x)=1+sec^2(x),

OpenStudy (anonymous):

because then you'd be moving the one to another side of the equation when you should really subtract it on both sides so yeah you'd have tan^2(x)=sec^2(x)-1

OpenStudy (anonymous):

so this is how I arranged it: 1-sec^2(x)--> 1-(tan^2(x)+1)-->(1-tan^2(x)-1)--->-tan^2(x). then I did (cos(x)/tan(x)) * (-tan^2(x)) --->cos(x) * -tan^2(x)---> cos(x)* (sin^2(x))/cos^2(x)) --> -sinx^2(x)/cosx that's what i have so far, am I on the right track?

OpenStudy (mertsj):

What is this equal to? Or are you just supposed to simplify it?

OpenStudy (anonymous):

simplify!

OpenStudy (anonymous):

I got a new result of -sin(x)

OpenStudy (anonymous):

is that okay?

OpenStudy (mertsj):

\[\frac{\cos x}{\frac{\sin x}{\cos x}}(1-\frac{1}{\cos ^2x})=\frac{\cos ^2x}{\sin x}(1-\frac{1}{\cos ^2x})\]

OpenStudy (anonymous):

is that your final result?

OpenStudy (mertsj):

no

OpenStudy (mertsj):

\[\frac{\cos ^2x}{\sin x}(\frac{\cos ^2x-1}{\cos ^2x})=\frac{-\sin ^2x}{\sin x}=-\sin x\]

OpenStudy (anonymous):

is it alright if I did 1-sec^2(x)--> 1-(tan^2(x)+1)?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thank you!!

OpenStudy (mertsj):

yw